python - Write a program that asks the user to enter how much their dinner bill was. Then ask them how many people were in their party. If there were more than 6 people in their party automatically calculate a 20% tip. If there were 6 or less people in their party ask them what percentage they would like to tip. In all cases, print out the total amount of the tip at the end of the program.
Answer #1
bill = float(input("Enter dinner bill: "))
n = int(input("How many people:"))
tip = 0
if(n>6):
    tip = bill*(20/100)
else:
    percent = int(input("How much percent you want to give: "))
    tip = bill * (percent / 100)
total = bill + tip
print("Total amount =",total)